A current of 2.5 A flowing in a solenoid having
self-inductance 3.8 H is reduced steadily to
zero in 1.0 ms. Calculate the magnitude of the
back emf of the solenoid while the current is
being switched off.
1
Expert's answer
2019-05-06T10:33:56-0400

Solution. The magnitude of the induced emf is equal to


"\Epsilon=\frac {\Delta \Phi} {\Delta t}"

where


"\Delta \Phi"

 is change in flux; t is time.

On the other hand the magnitude of the change in flux can be represented as


"\Delta \Phi=L \Delta I"

where L is inductance; I is current. Combining the two formulas we get


"\Epsilon=\frac {L \Delta I} {\Delta t}."

According to the data in the condition of the problem we get


"\Epsilon=\frac {3.8*2.5} {10^{-3}}=9500V"

Answer. 9500V